OpenStudy (anonymous):

Simplify the expression and rationalize the denominator \[\frac{1}{sqrt{t}}(\frac{1}{sqrt{t}}-1)\]

OpenStudy (anonymous):

\[\frac{1}{\sqrt{t}}(\frac{1}{\sqrt{t}}-1)\]

OpenStudy (anonymous):

The answer is... \[\frac{1-\sqrt{t}}{t}\]

OpenStudy (callisto):

\[\frac{1}{\sqrt{t}}(\frac{1}{\sqrt{t}}-1)\]\[=\frac{1}{\sqrt{t}}(\frac{1}{\sqrt{t}}-\frac{\sqrt{t}}{\sqrt{t}})\]\[=\frac{1}{\sqrt{t}}(\frac{1-\sqrt{t}}{\sqrt{t}})\]\[=\frac{1-\sqrt{t}}{\sqrt{t}(\sqrt{t})}\]\[=...?!\]

OpenStudy (anonymous):

But I got... \[\frac{t-t \sqrt{t}}{t}\]

OpenStudy (callisto):

Can you show your steps?

OpenStudy (anonymous):

I multiplied the coefficient in front first instead of doing what you did inside first: \[\frac{1}{t}-\frac{1}{\sqrt{t}}\] \[\frac{\sqrt{t}-t}{t\sqrt{t}}\] \[\frac{\sqrt{t}-t}{\sqrt[3]{t^2}}*\frac{\sqrt[3]{t^2}}{\sqrt[3]{t^2}}\] And then the answer I got...

OpenStudy (callisto):

\[t\sqrt{t} = \sqrt[2]{t^3}\]

OpenStudy (anonymous):

Woops yeah but now my answer is even mre wrong...\[\frac{\sqrt{t}-t}{t^2}\]

OpenStudy (callisto):

Let me try to do it using your method. \[\frac{1}{\sqrt{t}}(\frac{1}{\sqrt{t}}-1)\]\[=\frac{1}{t}-\frac{1}{\sqrt{t}}\]\[=\frac{\sqrt{t}}{t\sqrt{t}}-\frac{t}{t\sqrt{t}}\]\[=\frac{\sqrt{t}-t}{t\sqrt{t}}\]\[=\frac{\sqrt{t}(1-\sqrt{t})}{t\sqrt{t}}\]\[=\frac{1-\sqrt{t}}{t}\]

OpenStudy (anonymous):

\[= \frac{1}{t} - \frac{1}{\sqrt{t}} = \frac{1 - \sqrt{t}}{t}\]

OpenStudy (anonymous):

I don't understand after your second to last line

OpenStudy (callisto):

\[=\frac{\sqrt{t}}{t\sqrt{t}}-\frac{t}{t\sqrt{t}}\] This?

OpenStudy (anonymous):

I think there is no need of rationalizing the denominator by LCM we can make the denominator rational.. See. first distribute 1/root(t).. Then take the LCM you have your answer..

OpenStudy (callisto):

@waterineyes You're skipping too many steps.

OpenStudy (anonymous):

What what is the need of doing that steps?? First distribute: \[\frac{1}{t} - \frac{1}{\sqrt{t}}\] Now the LCM is t and solve further..

OpenStudy (callisto):

If the asker understands, that's fine.

OpenStudy (anonymous):

If the asker does not understand then I will explain it more..

OpenStudy (callisto):

So, now, it's time to check if @purplec16 understands.

OpenStudy (anonymous):

I am a visual learner. And I don't understand how you rationalize the denominator and where you got the square of t to take out from the numerator

OpenStudy (anonymous):

Can I explain you what Callisto did??

OpenStudy (anonymous):

Sure :D

OpenStudy (anonymous):

Tell me callisto's which step you are not getting??

OpenStudy (anonymous):

The last two

OpenStudy (anonymous):

In which reply of Callisto, there are these two steps??

OpenStudy (anonymous):

Let me try to do it using your method. In this reply or his earlier own reply??

OpenStudy (callisto):

Before \[=\frac{\sqrt{t}(1-\sqrt{t})}{t\sqrt{t}}\]add a step, so it becomes \[=\frac{\sqrt{t}-(\sqrt{t})(\sqrt{t})}{t\sqrt{t}}\]\[=\frac{\sqrt{t}(1-\sqrt{t})}{t\sqrt{t}}\]

OpenStudy (callisto):

It's taking out the common factor.

OpenStudy (anonymous):

I see it now... but how do you rationalize...

OpenStudy (anonymous):

See, callisto has taken root(t) common there and it will cancel with root(t) in the denominator..

OpenStudy (callisto):

I cross out the common factor. root(t)

OpenStudy (anonymous):

You have problem in here in rationalising or you don't know about rationalising??

OpenStudy (anonymous):

o.O...

OpenStudy (anonymous):

I know..

OpenStudy (callisto):

\[=\frac{\sqrt{t}(1-\sqrt{t})}{t\sqrt{t}}\]\[=\frac{\cancel{\sqrt{t}}(1-\sqrt{t})}{t\cancel{\sqrt{t}}}\]\[=\frac{(1-\sqrt{t})}{t}\]

OpenStudy (anonymous):

Cool... Thanks so much both!!!... my way is harder I think I will try to remember to use yours.

OpenStudy (callisto):

Not to remember, but to understand. When you understand, you can use it easier. When you remember, you may forget.

OpenStudy (anonymous):